# Injective flows for star-like manifolds

### v2

@@ A.6.3 Penalized likelihood regression
A.6.4 Bayesian mixing model
In this section we provide details about the experiments concerning Bayesian mixing models.
@@ A.6.4 Bayesian mixing model
We compare to a Metropolis Hastings sampler that can handle mixtures of Gaussians targets well in a non-sparse setting. It uses proposals that are by construction on the manifold by sampling from a downscaled Dirichlet distribution centered on the current state. We tried different proposals but did not observe any relevant improvement. As we showed, this sampler however struggles when significant sparseness is present. Our sampler uses 1000 chains each with 100’000 samples. In contrast, the proposed injective flow did not require any fine-tuning and learnt sparse solutions. For a fair comparison, we set the runtime of the flow to match that of the MCMC samplers and roughly similar memory footprint (±$\pm 20\%$) so both methods are compared at similar computational costs. Lastly, we checked that running the sampler for a significantly longer time (5×5$\times$5 ×longer) did not achieve the level of performance of the flow.

### v3

@@ A.6.3 Penalized likelihood regression
A.7 Approximate vs exact Jacobian: sample quality and runtime
Figure 10: Additional comparison between two identical models trained with the exact Jacobian, as we propose, according to Eq. (4) and with the Hutchinson trace estimator, commonly used in injective papers. The experimental setup is described in Section 5.1. We compare samples of the two models with samples from the true distribution as a function of the dimension and of the number of samples used in the Hutchinson trace estimator ($n=1,10,50,100$). Note that the number of samples is upper bounded by the dimensionality of the problem. We quantify the similarity of the samples with the samples from the true distribution with (a) the Energy distance MMD (“energy”) and (b) the Sinkhorn divergence (“sinkhorn”). The Energy distance MMD is computed using the kernel $-\|x-y\|_{2}$. The Sinkhorn divergence interpolates between Wasserstein (blur=0) and kernel (blur=$\infty$) distances and we used the default value blur=$0.05$. Below (c) we show the runtime of our exact model compared to the approximate one as a function of increasing number of samples.
A.7.1 Bayesian mixing model
@@
We compare to a Metropolis Hastings sampler that can handle mixtures of Gaussians targets well in a non-sparse setting. It uses proposals that are by construction on the manifold by sampling from a downscaled Dirichlet distribution centered on the current state. Despite trying different proposals we did not observe any relevant improvement. As shown, this sampler however struggles when significant sparseness is present. Our sampler uses 1000 chains each with 100’000 samples. In contrast, the proposed injective flow did not require any fine-tuning and learnt sparse solutions. For a fair comparison, we set the runtime of the flow to match that of the MCMC samplers and used roughly similar memory footprint (±$\pm 20\%$). Lastly, we checked that running the sampler for a significantly longer time (5×5$\times$5 ×) did not achieve the performance of the flow.